OpenStudy (anonymous):

What is the value of b? –7b = –42 b =

OpenStudy (anonymous):

divde or subtrat 42 with 7

OpenStudy (jacobciezki):

You times

OpenStudy (anonymous):

so 42 times 7

OpenStudy (jacobciezki):

No -7 *6=-42

OpenStudy (jacobciezki):

Try that

OpenStudy (anonymous):

brb got to take ashley to school

OpenStudy (jacobciezki):

Ok 👌 who that

OpenStudy (anonymous):

my sister

OpenStudy (jacobciezki):

Ok 👌 PM me when you get bavk

OpenStudy (texaschic101):

divide both sides by -7

OpenStudy (jacobciezki):

B is 6

OpenStudy (anonymous):

B is 6. Verification : -7*6 = -42. So yea :P

OpenStudy (texaschic101):

correct ^^ -7b = -42 --- divide both sides by -7 b = -42/-7 --- neg divided by a neg is a pos b = 6

OpenStudy (jacobciezki):

See I knew it

OpenStudy (texaschic101):

yes you did :)

OpenStudy (anonymous):

It is pretty easy, lol.

OpenStudy (jacobciezki):

Yea

OpenStudy (texaschic101):

lol...yes

OpenStudy (jacobciezki):

Ohhh lol girls are so smart now lol

OpenStudy (anonymous):

backkkkkkkk

OpenStudy (jacobciezki):

Ok 👌 😙

OpenStudy (anonymous):

262

OpenStudy (jacobciezki):

What

OpenStudy (anonymous):

if you times 7 and 6 you get 262

OpenStudy (jacobciezki):

No times -7*6=-42

OpenStudy (anonymous):

sorry;-; my cal

OpenStudy (jacobciezki):

Lol